The price of the used car is $23,500. Sales tax on this car is 7% of the price of the vehicle. You intend to finance the entire cost of the car and sales tax, less a down payment of $1,200. You intend to finance the car for 48 months and your car payment will be $551.44 per month.

Your answer should include the following information

* Calculate and discuss the amount financed, the installment price of the new car, and the finance charges you will pay over the life of the loan.
Answer #1

Solution:

Amount financed = (Cost + sales tax) - Down payment = ($23,500 + $23,500*7%) - $1,200 = $23,945

Installment price of new car = Installment + Down payment = ($551.44*48) + $1,200 = $27,669.12

Finance charges to be paid over the life of the loan = $27,669.12 - ($23,500*107%) = $2,524.12